OpenStudy (anonymous):

find y' dy where y = 2/x^3+x -- dx x=1

OpenStudy (chaise):

y = 2/x^3+x dy/dx = 2/3x^2+1 x = 1 (2/1^3)+1 = 3

myininaya (myininaya):

\[y=\frac{2}{x^3}+x?\]

OpenStudy (anonymous):

yes and dy/dx x=1

myininaya (myininaya):

\[y=2x^{-3}+x => y'=2(-3)x^{-3-1}+1=-6x^{-4}+1=\frac{-6}{x^4}+1\]

myininaya (myininaya):

\[y'|_{x=1}=\frac{-6}{(1)^4}+1=-6+1=-5\]

OpenStudy (anonymous):

awesome I also got -5 but was not sure if I did it right

myininaya (myininaya):

great!

OpenStudy (anonymous):

thx so much

myininaya (myininaya):

np :)